# Security and Real-time FPGA integration for Learned Image Compression

### v1

@@ IV-E On-board experiments
TABLE XI: Example model (BPP=0.2) showing changes in PSNR and MS-SSIM rates at different steps of the workflow evaluated on the CLIC dataset
HD FHD UHD
PSNR (dB) MS-SSIM (dB) PSNR (dB) MS-SSIM (dB) PSNR (dB) MS-SSIM (dB)
Float-Model 33.95 15.71 33.83 15.12 32.91 14.64
Pruning 32.74 15.37 32.15 14.44 32.43 14.22
QAT+WM 32.61 15.32 31.89 14.58 31.71 14.17
Next, we deploy the pruned, quantized and watermarked xmodels on the target boards for the final evaluation on the ZCU102 and KV260 FPGAs (Fig.7 shows the setup). Fig.8 and 9 report the RD curves on the Kodak and CLIC pictures that were left out of the training set. The figures show that our proposed FPGA compression platform outperforms the standardized codecs in most instances.
Table.XI shows the changes in RD efficiency from the original floating point model we trained to the final model deployed on the hardware. The platform utilizes the target boards’ embedded processors for pre- and post-processing, introducing software overhead, particularly in patch-based inference. This overhead is accounted for in the end-to-end latency.
@@
Figure 9: PSNR (left) and MS-SSIM (right) rate curves on the CLIC dataset
IV-F System-Level Performance and Hardware Deployment Complexity Analysis
@@
Table.X reports the system-level features in terms of Frames Per Second (FPS) for the hardware deployed model and the end-to-end pipeline. The combination of pruning and quantization enables significant FPS boots compared to similar works. Compared to [15], using the ZCU102 which has more resources than the KV260, the proposed platform achieves a speedup of 9.2×, 8.2× and 7.6× for the HD, FHD and UHD models respectively. A slight speedup is achieved compared to [16] if we assume their results are end-to-end and include software overhead. The FPS difference between UHD and FHD resolutions is primarily attributed to optimized parallelism in the FPGA’s inference process. Despite UHD images containing four times the number of pixels compared to FHD, the FPGA architecture efficiently balances the increased pixel count with throughput optimization, effectively utilizing more of the available resources which results in higher DSP efficiency. Additionally, the system benefits from Vitis AI’s optimized layer execution and pipelined memory access, enabling it to maintain high FPS despite the increased computational load.
@@
Table.VII shows the resource utilization for both boards and the work in [15], our ZCU102 is more resource efficient as it uses 3 DPU cores for the HD LIC task. Table.VI shows the power consumption on the ZCU102 board as measured placing shunt resistors on each of the power rails that interface with the FPGA. Each shunt resistor is linked to a TI-INA226 device, capable of measuring both voltage and current and connected to an I2C bus for convenient data collection and real-time voltage and current monitoring. We employ an embedded application running on the ARM processor to retrieve the relevant current inputs and voltage attributes, giving us PS, PL, and total power. For the KV260, we use Xilinx’s font-end platform management utility—xmutil, which allows for real-time monitoring of the embedded environment.Compared to [21], our model achieves a 70% reduction in power consumption for UHD resolution. Unlike [22], which reports results excluding software overhead, we include end-to-end performance metrics for comprehensive evaluation. Finally, Table.VI compares the proposed platform energy footprint with the related works and a GPU. Unsurprisingly, FPGAs outperform the GPU in the per-frame energy consumption; compared to [15] and [16], our platform exhibits a better energy budget for both end-to-end compression and decompression tasks.
IV-G Overhead of DRM Scheme
@@
The results from Tables VIII and IX, demonstrate that the introduction of Quantization Aware Watermarking and decryption (proposed DRM Scheme) into the FPGA accelerated Learned Image Compression pipeline for both KV260 and ZCU102 has a minimal impact on FPS and power consumption. For both platforms, the FPS values remain largely unaffected, with only minor reductions observed, indicating that the watermarking and decryption processes do not significantly degrade real-time performance. The small decrease in FPS, a -1.5% for HD on KV260 and -2.1% for HD on ZCU102, reflects the added computational overhead of DRM but does not hinder the devices’ ability to handle real-time image compression tasks. Power consumption also shows a modest increase, with +4.0% for HD and similar values for FHD and UHD, which is expected due to the additional processing required for watermark extraction and decryption. However, this increase remains within a reasonable range and should not pose a major concern in scenarios where security is prioritized. The Baseline power consumption is slightly lower than that of the Baseline + DRM scheme, as expected, but the overall power overhead is minimal. These results suggest that DRM can be effectively integrated into FPGA-based systems for secure image compression without significantly compromising performance or energy efficiency. The slight penalties in FPS and power consumption are outweighed by the enhanced security provided by DRM, making these devices suitable for applications that require secure image compression without sacrificing real-time performance.

### v2

@@ IV-D On-board experiments
Next, we deploy the pruned, quantized, and watermarked x-models onto the target FPGA boards for final evaluation on the ZCU102 and KV260 platforms. The hardware setup is shown in Fig. 8. The rate-distortion (RD) performance of the deployed models is analyzed on unseen images from the Kodak and CLIC datasets, as shown in Fig.9 and Fig.10, To facilitate comparison, MS-SSIM values are converted to dB, providing a more intuitive visualization of the relative quality differences between models.
These results demonstrate that our FPGA-based compression platform achieves superior performance compared to standardized codecs across most scenarios. Furthermore, Table X quantifies the impact of the full model optimization pipeline, from the original floating-point model to the final deployed hardware implementation.
The platform utilizes the target boards’ embedded processors for pre- and post-processing, introducing software overhead, particularly in patch-based inference. This overhead is accounted for in the end-to-end latency.
@@
Figure 10: PSNR (left) and MS-SSIM (right) rate curves on the CLIC2020 dataset
IV-E System-Level Performance and Hardware Deployment Complexity Analysis
@@
Table.XIII reports the system-level features in terms of Frames Per Second (FPS) for the hardware deployed model and the end-to-end pipeline. The combination of pruning and quantization enables significant FPS boots compared to similar works. Compared to [19], using the ZCU102 which has more resources than the KV260, the proposed platform achieves a speedup of 9.2×, 8.2× and 7.6× for the HD, FHD and UHD models respectively. A slight speedup is achieved compared to [20] if we assume their results are end-to-end and include software overhead. The FPS difference between UHD and FHD resolutions is primarily attributed to optimized parallelism in the FPGA’s inference process. Despite UHD images containing four times the number of pixels compared to FHD, the FPGA architecture efficiently balances the increased pixel count with throughput optimization, effectively utilizing more of the available resources which results in higher DSP efficiency. Additionally, the system benefits from Vitis AI’s optimized layer execution and pipelined memory access, enabling it to maintain high FPS despite the increased computational load.
@@
Table.XII shows the resource utilization for both boards and the work in [19], our ZCU102 is more resource efficient as it uses 3 DPU cores for the HD LIC task. Table.XI shows the power consumption on the ZCU102 board as measured placing shunt resistors on each of the power rails that interface with the FPGA. Each shunt resistor is linked to a TI-INA226 device, capable of measuring both voltage and current and connected to an I2C bus for convenient data collection and real-time voltage and current monitoring. We employ an embedded application running on the ARM processor to retrieve the relevant current inputs and voltage attributes, giving us PS, PL, and total power. For the KV260, we use Xilinx’s font-end platform management utility—xmutil, which allows for real-time monitoring of the embedded environment.Compared to [21], our model achieves a 70% reduction in power consumption for UHD resolution. Unlike [22], which reports results excluding software overhead, we include end-to-end performance metrics for comprehensive evaluation. Finally, Table.XI compares the proposed platform energy footprint with the related works and a GPU. Unsurprisingly, FPGAs outperform the GPU in the per-frame energy consumption; compared to [19] and [20], our platform exhibits a better energy budget for both end-to-end compression and decompression tasks.
IV-F Overhead of DRM Scheme
TABLE XIV: Impact of DRM on FPS and Energy (J/frame) for ZCU102. DRM refers to the combination of watermark-extraction and decryption applied to the baseline model.
Resolution Baseline Baseline + DRM $\Delta$ Energy (%) $\Delta$ FPS (%)
FPS Power [J/frame] FPS Power [J/frame]
HD (1280x720) 62.5 2.80 61.2 2.88 +2.9% -2.1%
FHD (1920x1080) 25.5 5.30 24.2 5.44 +2.6% -5.1%
UHD (3840x2160) 15.8 19.80 14.3 20.23 +2.2% -9.5%
TABLE XV: Impact of DRM on FPS and Energy (J/frame) for KV260. DRM refers to the combination of watermark-extraction and decryption applied to the baseline model.
Resolution Baseline Baseline + DRM $\Delta$ Energy (%) $\Delta$ FPS (%)
FPS Power [J/frame] FPS Power [J/frame]
HD (1280x720) 26.8 3.00 26.4 3.12 +4.0% -1.5%
FHD (1920x1080) 9.9 8.40 9.2 8.59 +2.3% -7.1%
UHD (3840x2160) 4.8 31.50 4.2 31.71 +0.7% -12.5%
@@ IV-F Overhead of DRM Scheme
The results from Tables XV and XIV, demonstrate that the introduction of Quantization Aware Watermarking and decryption (proposed DRM Scheme) into the FPGA accelerated Learned Image Compression pipeline for both KV260 and ZCU102 has a minimal impact on FPS and power consumption. For both platforms, the FPS values remain largely unaffected, with only minor reductions observed, indicating that the watermarking and decryption processes do not significantly degrade real-time performance. The small decrease in FPS, a -1.5% for HD on KV260 and -2.1% for HD on ZCU102, reflects the added computational overhead of DRM but does not hinder the devices’ ability to handle real-time image compression tasks. Power consumption also shows a modest increase, with +4.0% for HD and similar values for FHD and UHD, which is expected due to the additional processing required for watermark extraction and decryption. However, this increase remains within a reasonable range and should not pose a major concern in scenarios where security is prioritized. The Baseline power consumption is slightly lower than that of the Baseline + DRM scheme, as expected, but the overall power overhead is minimal. These results suggest that DRM can be effectively integrated into FPGA-based systems for secure image compression without significantly compromising performance or energy efficiency. The slight penalties in FPS and power consumption are outweighed by the enhanced security provided by DRM, making these devices suitable for applications that require secure image compression without sacrificing real-time performance.